OpenStudy (anonymous):

write an equation of a circle. area 4pi, center (1, -2)

OpenStudy (anonymous):

2^2=(y+2)^2+(x-1)^2 I think

OpenStudy (anonymous):

\[11.3*10^3/2.7*10^3)^(1/3)=ra/ri\]

OpenStudy (anonymous):

The more important thin is that you know how to do it Kayla. Did I get it right? if so do you know what I did to get answer?

OpenStudy (anonymous):

Yea because 2^2, but it has pi, so im checking to see if you do anything with the pi...